OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (bradely):

(y-y1)=m(x-x1) (y+2)=(2/5)(x-5) 5(y+2)=2(x-5) 5y+10=2x-10 2x-5y-20=0 Source: http://www.mathskey.com/question2answer/